OpenStudy (anonymous):

solve for the specified variable 18-6a=42-4(a+3)

OpenStudy (anonymous):

distribute the -4 to everything in the parentheses

OpenStudy (anonymous):

what will that look like i am lost

OpenStudy (anonymous):

18-6a=42-4a-12

OpenStudy (anonymous):

then you add like terms on whats on the right side. so you add 42 and -12 and get 30

OpenStudy (anonymous):

18-6a=30-4a is the equation you have now. so you need to move the variable a so they are on the same side.

OpenStudy (anonymous):

18-6a+-4(a+3)

OpenStudy (anonymous):

that one was worng one

OpenStudy (anonymous):

wait so you need to solve for 18-6a+-4(a+3) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you still distribute the -4

OpenStudy (anonymous):

Listen to @xlilly because to me this is hard sorry

OpenStudy (anonymous):

ok so distribute the -4 with the +3?

OpenStudy (anonymous):

18-6a+-4a-12 is what you should get after you distribute

OpenStudy (anonymous):

to everything that is in the parentheses

OpenStudy (anonymous):

ok i got it so far

OpenStudy (anonymous):